OpenStudy (anonymous):

find the smallest possible integer value of n for which 99n is a multiple of 24.

OpenStudy (nurali):

For a number to be a multiple of 24, it must contain all the factors of 24; which are 2 x 2 x 2 x 3. If you break down 99 into prime factors, it is 3 x 3 x 11. So of (2,2,2,3), 99 contains the 3 but does not have any of the 2s. So the smallest number would be 3 x 3 x 11 x (2 x 2 x 2) = 792 and n = 8.

OpenStudy (anonymous):

another approach: 99n/24 = (99/24 )*n = (4 + 1/8)*n it is obvious now that it is 8.

OpenStudy (nurali):

Welcome to Openstudy @bushraallahyar